If cos theta = -3/5, fund cos 2 theta and express in simplest form.
Answer by stanbon(75887)

: If cos theta = -3/5, fund cos 2 theta and express in simplest form.
---------------
cos(2t) = cos^2(t) - sin^2(t) = 2cos^2(t)-1 = 2(9/25)-1 = -7/25
---------------------
Cheers,
Stan H.
------------------